OpenStudy (anonymous):

Simplify, square root of 75 over 256

OpenStudy (anonymous):

\[\sqrt{75\256}\] = \[5\sqrt{2}\]/\[8\sqrt{2}\] = 5/8

OpenStudy (anonymous):

thanks:)))